# Hao Ma1, Zhiyuan Peng2, Xu Li3, Yukai Li1, Mingjie Shao1, Qiuqiang Kong4, and Ju Liu1

### v1

@@ Abstract
Language-queried target sound extraction (TSE) aims to extract specific sounds from mixtures based on language queries. Traditional fully-supervised training schemes require extensively annotated parallel audio-text data, which are labor-intensive. We introduce a language-free training scheme, requiring only unlabelled audio clips for TSE model training by utilizing the multi-modal representation alignment nature of the contrastive language-audio pre-trained model (CLAP). In a vanilla language-free training stage, target audio is encoded using the pre-trained CLAP audio encoder to form a condition embedding for the TSE model, while during inference, user language queries are encoded by CLAP text encoder. This straightforward approach faces challenges due to the modality gap between training and inference queries and information leakage from direct exposure to target audio during training. To address this, we propose a retrieval-augmented strategy. Specifically, we create an embedding cache using audio captions generated by a large language model (LLM). During training, target audio embeddings retrieve text embeddings from this cache to use as condition embeddings, ensuring consistent modalities between training and inference and eliminating information leakage. Extensive experiment results show that our retrieval-augmented approach achieves consistent and notable performance improvements over existing state-of-the-art with better generalizability.
@@ I Introduction
Figure 1: Overview of our proposed retrieval-augmented language-free training scheme for language-queried target sound extraction.
@@
Humans possess a remarkable ability to focus on specific sounds in noisy environments, a phenomenon known as the cocktail party effect. In signal processing, sound separation [1, 2] has been extensively researched to address this challenge. Early work focused on domain-specific signals like speech [3, 4] or music [5, 6], but recent advances in deep learning have led to universal sound separation (USS) [7], with the ambition to generalize the sound separation model to all kinds of sounds. A key development in USS is framing it as language-queried target sound extraction (TSE) [8, 9, 10, 11], where natural language descriptions are used to extract specific sounds from audio mixtures.
@@
Despite significant progress in language-queried TSE, practical challenges persist. Current TSE systems [8, 9, 11] heavily depend on large-scale text annotations of sound events for training. This fully-supervised approach requires extensive parallel audio-text datasets to enhance model accuracy and generalizability. However, the high labor cost of annotations limits the availability of parallel data, making it insufficient for scalable universal sound extraction. This scarcity, though under-explored in TSE, is a known challenge in other cross-modal tasks like text-to-image [12], text-to-audio [13], and audio captioning [14]. These studies [12, 13, 14] address the issue by leveraging modality-aligned embedding spaces from contrastive cross-modal pre-trained models, eliminating the need for extensive parallel data.
@@
Building on this approach, the contrastive language-audio pre-trained model (CLAP) [15] can be used to eliminate the need for text annotations in TSE. Typically, TSE models are trained using a target audio-text pair mixed with non-target audio, learning to extract the target sound in the audio mixture based on the text query. To bypass text annotations, CLAP can be cascaded with the TSE model. During training, the target audio is encoded into an audio embedding by the CLAP audio encoder, which the TSE model uses to extract the target sound. During testing, a text query is encoded by the CLAP text encoder into a text embedding for sound extraction. However, this vanilla approach assumes perfect alignment between text and audio embeddings, which is unrealistic. Two major challenges arise: the persistent modality gap between text and audio, worsened by training-testing inconsistencies, and the risk of overfitting due to the exposure of rich auditory details in the target audio embedding during training.
@@
To address these challenges, we propose a retrieval-augmented, language-free training paradigm for language-queried TSE models as detailed in Fig. 1. Before training, we use a large language model to generate diverse audio captions, which are encoded into text embeddings by the pre-trained CLAP text encoder and stored in an embedding cache. During training, we retrieve the most similar text embeddings based on the target audio embedding and use them as the condition embedding, guiding the TSE model in sound extraction. This retrieval approach mitigates the modality mismatch and information leakage issues. Additionally, we apply noise injection [12, 16, 14] to the conditional embedding as an augmentation strategy, enhancing model generalizability. Extensive experiments show that our retrieval-augmented approach significantly improves performance and generalizability compared to existing state-of-the-art methods. By relaxing the need for parallel data, our method scales easily for large-scale training and outperforms previous supervised training schemes across multiple benchmarks, achieving a 1-2 dB improvement in signal-to-distortion ratio (SDR).
@@ II Language-queried Target Sound Extraction
In this section, we propose to address the challenging task of language-queried target sound extraction. We first present the traditional fully-supervised training scheme when parallel audio-text data is available. Then, we introduce the proposed language-free training scheme with no need for parallel audio-text data.
@@ II-A Fully-Supervised Training with Parallel Data
In fully-supervised training, we have access to audio-text pairs ${\bm{d}}=\{{\bm{x}},{\bm{y}}\}$, where ${\bm{x}}$ is an audio clip and ${\bm{y}}$ is its corresponding annotation text. In the mix-and-separate training pipeline, the audio mixture is constructed by sampling several audio clips, treating one of them as the target source and the others as noise, and then summing them as:
@@
where $\hat{{\bm{x}}}\in{\mathbb{R}}^{N}$ represents the target sound source prediction and ${\bm{c}}\in{\mathbb{R}}^{D}$ denotes the $D$-dimensional condition embedding, which is acquired by encoding the annotation text of the target sound by the CLAP text encoder both in the training and inference stages, which is:
@@ II-B Proposed Approach
In language-free training without parallel data, we only have access to audio clips without text annotations. The critical challenge is how to construct the condition embedding ${\bm{c}}$ without access to text annotations? Our methods take advantage of the modality-aligned embedding space of contrastive language-audio pre-trained models. In the following part of this section, we first present a vanilla language-free training scheme and the widely adopted Gaussian noise injection as an augmentation strategy. We then propose a retrieval-augmented strategy to further mitigate specific challenges posed by existing methods to achieve improved performance.
II-B1 Vanilla Language-Free Training
In the vanilla language-free training, the condition embedding ${\bm{c}}$ is constructed by directly encoding the target audio ${\bm{x}}$ with the pre-trained CLAP audio encoder as:
@@ II-B1 Vanilla Language-Free Training
During inference, we use the pre-trained CLAP text encoder to encode the user language query to control the TSE model. This vanilla language-free training scheme works thanks to the modality-aligned embedding space generated by pre-trained CLAP, but it also suffers from two performance degradation points: one is the modality gap arose from the imperfect alignment of audio and text embeddings [17, 18]; the other is the information leakage by directly exposing the target audio during training that makes the model prone to overfitting.
@@ II-B2 Gaussian Noise Injection
Gaussian noise injection has been proven to be effective in bridging the modality gap in many other cross-modal tasks [12, 16, 14]. This method models the gap between the audio and text embeddings extracted by pre-trained CLAP encoders on parallel audio and text as Gaussian noise with mean 00 and variance $\epsilon$. Based on the above modeling, the Gaussian noise is used to perturb the condition embeddings extracted from the target audio during language-free training of a TSE model as:
@@ II-B3 Retrieval-Augmented Strategy
Although the modality gap is effectively addressed by Gaussian noise injection, the issue of information leakage remains a point of concern. We further propose a retrieval-augmented strategy to address both the modality gap and the information leakage problem simultaneously. As shown in Fig. 1, before the model training, we first generate massive audio captions by prompting a large language model following [19]. These audio captions are then encoded by the pre-trained CLAP text encoder to create an embedding cache ${\bm{E}}\in{\mathbb{R}}^{M\times D}$, where $M$ is the number of audio captions generated by LLM.
Algorithm 1 retrieval-augmented language-free training scheme.
@@
9:    Compute similarity: $\mathrm{sim}\leftarrow{\bm{q}}\cdot{\bm{E}}^{T}\in\mathbb{R}^{B\times M}$;
@@
TABLE I: Laguage-queried TSE performance evaluation. “bal.” denotes AudioSet balanced_train and “unbal.” denotes AudioSet unbalanced_train. The term “Vanilla” denotes target audio embeddings w/o any augmentation as condition embeddings for training and “NI” is noise injection for short. Except for LASS and AudioSep, all the methods listed in this table are built upon CLAPSep with different training strategies. †: reproduction in [11], ‡: “others” in [9] is comprised of AudioCaps, Clotho v2, WavCaps, VGGSound.
@@
In training, the target audio ${\bm{x}}$ is encoded by the CLAP audio encoder as a query to retrieve the most similar text embedding in the embedding cache ${\bm{E}}$. The retrieved text embedding is considered as a condition embedding ${\bm{c}}$ and further augmented by a Gaussian perturbation as:
@@
where the retrieval function is detailed in Alg. 1 and ${\bm{n}}_{1}\sim N(\mathbf{0},\epsilon_{1}{\bm{I}})\in{\mathbb{R}}^{D}$. We explore the setting of $\epsilon_{1}$ in Section III-C.
By retrieving pre-encoded text embeddings, we address the modality gap because the queries are in the same text modality in both the training and inference phases. Meanwhile, by using retrieved condition embeddings, the TSE model does not directly see the target audio embeddings during its training, further avoiding information leakage. A more detailed procedure for retrieval-augmented language-free training of a language-queried target sound extraction model is given in Alg. 1.
@@ III-A Datasets and Evaluation Metrics
Firstly, we perform quick preliminary experiments on the balanced_train split of AudioSet, which contains 20,550 audio clips, to evaluate the effectiveness of our proposed method and determine the hyperparameter settings. We then scale up the training data to include all collected audio clips from AudioSet and FreeSound to further validate the effectiveness of our method on large-scale non-parallel data. We refer to this large dataset as $\mathcal{DS}_{large}$.
We follow recipes in [9, 11] to prepare evaluation datasets. We perform evaluations on AudioCaps [22], Clotho v2 [23], the AudioSet test split, MUSIC21 [24] and ESC50 [25] to comprehensively validate the effectiveness of our proposed method. Note that in AudioCaps and Clotho v2, the annotations are human-written audio captions in natural language form, while in AudioSet, MUSIC21, and ESC50, the annotations are labels of sound events. In testing, we convert these labels to language-like audio captions by adding the prefix “The sound of ”.
@@
$\mathrm{SI\text{-}SDR}(\hat{{\bm{x}}},{\bm{x}})$ $=10*\mathrm{log}_{10}\left(\frac{||\frac{||\hat{{\bm{x}}}^{T}{\bm{x}}||^{2}}{||{\bm{x}}||^{2}}{\bm{x}}||^{2}}{||\frac{||\hat{{\bm{x}}}^{T}{\bm{x}}||^{2}}{||{\bm{x}}||^{2}}{\bm{x}}-\hat{{\bm{x}}}||^{2}}\right).$ (10)
@@ III-B Implementation Details
We build the language-free training framework upon the state-of-the-art (SOTA) language-queried target sound extraction model, CLAPSep [11]. For experiments on the balanced_train split, we train the model on a single RTX 3090 GPU with a batch size of 32 for a total of 96,450 steps. The initial learning rate is set to 1e-4, which decays exponentially by a factor of 0.32 at steps 31,150 and 64,300. For experiments on $\mathcal{DS}_{large}$, we train the model on four RTX 3090 GPUs with a global batch size of 64 for a total of 509,400 steps. The initial learning rate is set to 2e-4, which decays exponentially by a factor of 0.32 at steps 169,800 and 339,600. We adopt the negative SI-SDR between the model estimation and the target audio as an end-to-end loss function as detailed in equation 10.
@@ III-C Results and Analysis
Effectiveness of retrieval-augmented strategy. We show the performance of various language-free training schemes in Table I. As can be seen from these results, the vanilla strategy, i.e., the direct use of target audio embedding without any augmentation as the condition embedding for training, performs the worst. This indicates that the modality gap and information leakage introduced by this strategy cause a significant drop in the performance of the TSE model. The modality gap problem can be solved to some extent by Gaussian noise injection. As can be seen from the results in the table (Vanilla w/ NI on bal.), Gaussian noise injection brings more than 2dB performance improvement on average for the vanilla strategy. Further, the retrieval-augmented strategy proposed in this paper combined with a certain degree of noise injection (Retrieval w/ NI on bal.) performs optimally over all kinds of language-free training schemes. It is worth noting that the better performance achieved by our proposed method does not rely on Gaussian noise injection, which is shown by the performance gap between “Retrieval” and “Retrieval w/ NI” is smaller than that between vanilla strategies. This indicates that the proposed method has effectively addressed the modal gap and information leakage problems by retrieving pre-encoded text embeddings.

### v3

# Hao Ma1, Zhiyuan Peng2, Xu Li3, Yukai Li1, Mingjie Shao4, Qiuqiang Kong5, and Ju Liu1

@@ Abstract
Language-queried target sound extraction (TSE) aims to extract specific sounds from mixtures based on language queries. Traditional fully-supervised training schemes require extensively annotated parallel audio-text data, which are labor-intensive. We introduce a parallel-data-free training scheme, requiring only unlabelled audio clips for TSE model training by utilizing the contrastive language-audio pre-trained model (CLAP). In a vanilla parallel-data-free training stage, target audio is encoded using the pre-trained CLAP audio encoder to form a condition embedding, while during testing, user language queries are encoded by CLAP text encoder as the condition embedding. This vanilla approach assumes perfect alignment between text and audio embeddings, which is unrealistic. Two major challenges arise from training-testing mismatch: the persistent modality gap between text and audio and the risk of overfitting due to the exposure of rich acoustic details in target audio embedding during training. To address this, we propose a retrieval-augmented strategy. Specifically, we create an embedding cache using audio captions generated by a large language model (LLM). During training, target audio embeddings retrieve text embeddings from this cache to use as condition embeddings, ensuring consistent modalities between training and testing and eliminating information leakage. Extensive experiment results show that our retrieval-augmented approach achieves consistent and notable performance improvements over existing state-of-the-art with better generalizability.
@@ I Introduction
Humans possess a remarkable ability to focus on specific sounds in noisy environments, a phenomenon known as the cocktail party effect. In signal processing, sound separation [1, 2] has been extensively researched to address this challenge. Early work focused on domain-specific signals like speech [3, 4] or music [5, 6], but recent advances in deep learning have led to universal sound separation (USS) [7, 8], with the ambition to generalize to separate all kinds of sounds. A key development in USS is framing it as query-oriented target sound extraction (TSE) [9, 10, 11, 12, 13, 14, 15, 16, 17, 18], where auxiliary information such as language are introduced to specify what sound to extract.
@@
Despite significant progress in language-queried TSE, practical challenges persist. Current language-queried TSE systems [19, 16, 18] heavily depend on large-scale text annotations of sound events for training. This fully-supervised approach requires extensive parallel audio-text datasets to enhance model accuracy and generalizability. However, the high labor cost of annotations limits the availability of parallel data, making it insufficient for scalable universal sound extraction. This scarcity, though under-explored in TSE, is a known challenge in other cross-modal tasks like text-to-image [20], text-to-audio [21], and audio captioning [22]. These studies [20, 21, 22] address the issue by leveraging modality-aligned embedding spaces from contrastive cross-modal pre-trained models that are previously trained on massive cross-modal data pairs, eliminating the need for extensive parallel data for specific downstream tasks.
@@
Building on this paradigm, the contrastive language-audio pre-trained model (CLAP) [23] can be used to eliminate the need for text annotations in TSE. Typically, TSE models are trained using a target audio-text pair mixed with non-target audio, learning to extract the target sound in the audio mixture based on the text query. To bypass text annotations, CLAP can be cascaded with the TSE model. During training, the target audio is encoded into an audio embedding by the CLAP audio encoder, which the TSE model uses to extract the target sound. During testing, a text query is encoded by the CLAP text encoder into a text embedding for sound extraction. However, this vanilla approach assumes perfect alignment between text and audio embeddings, which is unrealistic. Two major challenges arise from training-testing mismatch: the persistent modality gap [24] between text and audio and the risk of overfitting due to the exposure of rich acoustic details in target audio embedding during training.
Figure 1: Overview of our proposed retrieval-augmented parallel-data-free training scheme for language-queried target sound extraction.
@@
To address these challenges, we propose a retrieval-augmented, parallel-data-free training paradigm for language-queried TSE models as detailed in Fig. 1. Before training, we use a large language model to generate diverse audio captions, which are encoded into text embeddings by the pre-trained CLAP text encoder and stored in an embedding cache. During training, we retrieve the most similar text embeddings based on the target audio embeddings and use them as the condition embeddings, guiding the TSE model in sound extraction. Using this retrieval-augmented strategy which retrieves pre-encoded text embedding for model training, we achieve direct alignment of query modalities between training and testing. Additionally, our approach effectively mitigates the leakage of fine-grained acoustic information present in target audio embeddings—information that query language cannot encapsulate during testing. We also investigate applying noise injection [20, 25, 22] to the condition embedding as an augmentation strategy, enhancing model generalizability. Extensive experiments show that our retrieval-augmented approach significantly improves performance and generalizability compared to existing methods. By relaxing the need for parallel data, our method scales easily for large-scale training and outperforms previous fully-supervised training schemes across multiple benchmarks, achieving a 1-2 dB improvement in signal-to-distortion ratio (SDR).
@@ II Language-queried Target Sound Extraction
In this section, we propose to address the challenging task of language-queried target sound extraction. We first present the traditional fully-supervised training scheme when parallel audio-text data is available. Then, we introduce the proposed parallel-data-free training scheme with no need for parallel audio-text data.
@@ II-A Fully-Supervised Training with Parallel Data
In fully-supervised training, we have access to audio-text pairs ${\bm{d}}=\{{\bm{x}},{\bm{y}}\}$, where ${\bm{x}}$ is an audio clip and ${\bm{y}}$ is its annotation text. In the mix-and-separate training pipeline, the audio mixture is constructed by sampling several audio clips, treating one of them as the target source and the others as noise, and then summing them as:
@@
where $\hat{{\bm{x}}}\in{\mathbb{R}}^{N}$ represents the predicted target sound source and ${\bm{c}}\in{\mathbb{R}}^{D}$ denotes the $D$-dimensional condition embedding, which is acquired by encoding the annotated text of the target sound by the CLAP text encoder both in the training and testing stages as:
@@ II-B Proposed Approach
In parallel-data-free training, we only have access to audio clips without text annotations. The critical challenge is how to construct the condition embedding ${\bm{c}}$ without access to text annotations? Our methods take advantage of the modality-aligned embedding space of contrastive language-audio pre-trained models. In the following part of this section, we first present a vanilla parallel-data-free training scheme and the widely adopted Gaussian noise injection as an augmentation strategy. We then propose a retrieval-augmented strategy to further mitigate specific challenges posed by existing methods to achieve improved performance.
II-B1 Vanilla Parallel-Data-Free Training
In vanilla parallel-data-free training, the condition embedding ${\bm{c}}$ is constructed by encoding the target audio ${\bm{x}}$ with the pre-trained CLAP audio encoder as:
@@ II-B1 Vanilla Parallel-Data-Free Training
During testing, we use the pre-trained CLAP text encoder to encode the user language query to control the TSE model. This vanilla strategy works thanks to the modality-aligned embedding space generated by pre-trained CLAP, but it also suffers from performance degradation mainly from the mismatch between training and testing: one is the modality gap arose from the imperfect alignment of audio and text embeddings [26, 24]; the other is the leakage of fine-grained acoustic details in target audio embedding, which cannot be captured by query language, making the model prone to overfitting.
@@ II-B2 Gaussian Noise Injection
Gaussian noise injection has been proven to be effective in bridging the modality gap in many other cross-modal tasks [20, 25, 22]. This method models the gap between the audio and text embeddings extracted by pre-trained CLAP encoders on parallel audio and text as Gaussian noise with mean 00 and variance $\epsilon$. Based on the above modeling, the Gaussian noise is used to perturb the condition embeddings extracted from the target audio during parallel-data-free training of a TSE model as:
@@ II-B3 Retrieval-Augmented Strategy
Algorithm 1 retrieval-augmented parallel-data-free training scheme.
@@
9:    Compute cosine similarity: $\mathrm{sim}\leftarrow\frac{{\bm{q}}{\bm{E}}^{T}}{||{\bm{q}}||||{\bm{E}}||}\in\mathbb{R}^{B\times M}$;
@@
TABLE I: Language-queried TSE performance evaluation. “bal.” denotes AudioSet balanced_train and “unbal.” denotes AudioSet unbalanced_train. The term “Vanilla” denotes target audio embeddings w/o any augmentation as condition embeddings for training and “NI” is noise injection for short. Except for LASS and AudioSep, all the methods listed in this table are built upon CLAPSep [18]. †: reproduction in [18], ‡: “others” in [16] is comprised of AudioCaps, Clotho v2, WavCaps, VGGSound.
@@
Although many previous works [20, 25, 22] have demonstrated that Gaussian noise injection effectively addresses the mismatch issue for cross-modal tasks, there still exists a performance gap compared to training with manually labeled data pairs [22]. To further bridge this gap, we introduce a retrieval-augmented strategy that directly aligns query modalities between training and testing while addressing the leakage of acoustic information that language-based queries cannot encapsulate. As shown in Fig. 1, before the model training, we first generate massive audio captions by prompting a large language model following [27]. These audio captions are then encoded by the pre-trained CLAP text encoder to create an embedding cache ${\bm{E}}\in{\mathbb{R}}^{M\times D}$, where $M$ is the number of audio captions generated by LLM. In training, the target audio ${\bm{x}}$ is encoded by the CLAP audio encoder as a query to retrieve the most similar text embedding in the embedding cache ${\bm{E}}$. The retrieved text embedding is considered as a condition embedding ${\bm{c}}$ and further augmented by a Gaussian perturbation as:
@@
where we retrieve the most similar text embedding in the embedding cache according to cosine similarity; and ${\bm{n}}_{1}\sim N(\mathbf{0},\epsilon_{1}{\bm{I}})\in{\mathbb{R}}^{D}$. We explore the setting of $\epsilon_{1}$ in Section III-C. A more detailed procedure for retrieval-augmented parallel-data-free training of a language-queried target sound extraction model is given in Alg. 1.
@@ III-A Datasets and Evaluation Metrics
We first perform quick preliminary experiments on the balanced_train split of AudioSet, which contains 20,550 audio clips, to evaluate the effectiveness of our proposed method and determine the hyperparameter settings. We then scale up the training data to include all collected audio clips from AudioSet and FreeSound to further validate the effectiveness of our method on large-scale non-parallel data. We refer to this large dataset as $\mathcal{DS}_{large}$.
We follow recipes in [16, 18] to prepare evaluation datasets. We perform evaluations on AudioCaps [30], Clotho v2 [31], the AudioSet test split, MUSIC21 [15] and ESC50 [32], which cover a wide range of natural sound and music to comprehensively validate the effectiveness of our proposed method. Note that in AudioCaps and Clotho v2, the annotations are human-written audio captions in natural language form, while in AudioSet, MUSIC21, and ESC50, the annotations are labels of sound events. In testing, we convert these labels to language-like audio captions by adding the prefix “The sound of ”.
@@
$\mathrm{SI\text{-}SDR}(\hat{{\bm{x}}},{\bm{x}})$ $=10*\mathrm{log}_{10}\left(\frac{||\frac{\hat{{\bm{x}}}^{\top}{\bm{x}}}{||{\bm{x}}||^{2}}{\bm{x}}||^{2}}{||\frac{\hat{{\bm{x}}}^{\top}{\bm{x}}}{||{\bm{x}}||^{2}}{\bm{x}}-\hat{{\bm{x}}}||^{2}}\right).$ (10)
@@ III-B Implementation Details
We build the parallel-data-free training framework upon the state-of-the-art (SOTA) language-queried target sound extraction model, CLAPSep [18], which is a transformer-based model. For experiments on the balanced_train split, we train the model on a single RTX 3090 GPU with a batch size of 32 for a total of 96,450 steps. The initial learning rate is set to 1e-4, which decays exponentially by a factor of 0.32 at steps 31,150 and 64,300. For experiments on $\mathcal{DS}_{large}$, we train the model on four RTX 3090 GPUs with a global batch size of 64 for a total of 509,400 steps. The initial learning rate is set to 2e-4, which decays exponentially by a factor of 0.32 at steps 169,800 and 339,600. We adopt the negative SI-SDR between the model estimation and the target audio as the loss function as detailed in equation 10.
@@ III-C1 Effectiveness of Retrieval-Augmented Strategy
We show the performance of various parallel-data-free training schemes in Table I. As can be seen from these results, the vanilla strategy, i.e., the direct use of target audio embedding without any augmentation as the condition embedding for training, performs the worst. This indicates that the mismatch during training and testing time introduced by this strategy causes a significant performance drop. The mismatch problem can be solved effectively by Gaussian noise injection. As can be seen from the results in the table (Vanilla w/ NI on bal.), Gaussian noise injection brings a performance improvement of more than 2dB on average for the vanilla strategy. Further, the retrieval-augmented strategy proposed in this paper combined with a certain degree of noise injection (Retrieval w/ NI on bal.) performs optimally over all kinds of parallel-data-free training schemes. It is worth noting that the better performance achieved by our proposed method does not rely on Gaussian noise injection, which is shown by the performance gap between “Retrieval” and “Retrieval w/ NI” is smaller than that between vanilla strategies. This indicates that the proposed method has more effectively addressed the training-testing mismatch problem by retrieving pre-encoded text embeddings.
III-C2 Parallel-Data-Free Versus Weakly-Supervised